OpenStudy (anonymous):

Solve using the substitution method. 8x – 5y = 3 7x + y = 51

OpenStudy (unklerhaukus):

so re shape the second equation to be y = ... \[y=51-7x\] now substitute this into your first equation for y

OpenStudy (unklerhaukus):

\[8x-5(51-7x)=3\] and solve for x

OpenStudy (anonymous):

i dont know what to do next

OpenStudy (unklerhaukus):

expand the bracket, ie -5 times 51 , and -5 times -7x

OpenStudy (anonymous):

8x-225-35x=3

OpenStudy (unklerhaukus):

careful now -5 times -7x is +35x

OpenStudy (anonymous):

the x needs to cancel next right?

OpenStudy (unklerhaukus):

yeah

OpenStudy (anonymous):

thanks

OpenStudy (unklerhaukus):

\[8x – 5y = 3\]\[8x-5(51-7x)=3\]\[8x-225+35x=3\] ...

OpenStudy (anonymous):

43x=222

OpenStudy (unklerhaukus):

check that again we are adding 225 to both sides

OpenStudy (anonymous):

8x-225+35x=3 +225 +225 cancel 8x+35x=228

OpenStudy (unklerhaukus):

that is good

OpenStudy (anonymous):

im not done tho right?

OpenStudy (unklerhaukus):

i am not getting a nice answer for this question

OpenStudy (unklerhaukus):

i keep getting x=228/43

OpenStudy (anonymous):

me too..that is why i asked for help bc i questioned myself

OpenStudy (unklerhaukus):

but i think ive made an error somewhere

OpenStudy (unklerhaukus):

x=6,y=9 is a working solution but im not sure how to get there, or what we did wrong

OpenStudy (anonymous):

i will turn in what we did and get sum feedback

OpenStudy (unklerhaukus):

5 x 51 = 255

OpenStudy (unklerhaukus):

\[8x-5y=3\qquad (i)\]\[7x+y=51\qquad(ii)\] \[(ii)\rightarrow\qquad y=51-7x\qquad(iii)\]\[(iii)\rightarrow(i)\qquad 8x-5(51-7x)=3\]\[\qquad 8x-255+35x=3\]\[\qquad\qquad 43x=258\]\[x=\frac{258}{43}\]\[x=6\qquad (iv)\] \[(iv)\rightarrow(ii)\qquad 7(6)+y=51\]\[42+y=51\]\[y=9\]